# Learning Perceptually Relevant Temporal Envelope Morphing

### v1

@@ Abstract
Temporal envelope morphing, the process of interpolating between the amplitude dynamics of two audio signals, is an emerging problem in generative audio systems that lacks sufficient perceptual grounding. Morphing of temporal envelopes in a perceptually intuitive manner should enable new methods for sound blending in creative media and for probing perceptual organization in psychoacoustics. However, existing audio morphing techniques often fail to produce intermediate temporal envelopes when input sounds have distinct temporal structures; many morphers effectively overlay both temporal structures, leading to perceptually unnatural results. In this paper, we introduce a novel workflow for learning envelope morphing with perceptual guidance: we first derive perceptually grounded morphing principles through human listening studies, then synthesize large-scale datasets encoding these principles, and finally train machine learning models to create perceptually intermediate morphs. Specifically, we present: (1) perceptual principles that guide envelope morphing, derived from our listening studies, (2) a supervised framework to learn these principles, (3) an autoencoder that learns to compress temporal envelope structures into latent representations, and (4) benchmarks for evaluating audio envelope morphs, using both synthetic and naturalistic data, and show that our approach outperforms existing methods in producing temporally intermediate morphs. All code, models, and datasets will be made publicly available upon publication.
@@ 1 Introduction
Figure 1: Method overview. In Stage 1, we repurpose an audio autoencoder to learn a latent representation of temporal envelopes. In Stage 2, we train a twin (order-invariant) mapper network (2b) in a supervised setup on synthetic training data that encodes perceptually-grounded morphing rules (2a). During training, the autoencoder is kept frozen while the mapper network learns to morph the envelope embeddings.
@@ 3 Perceptually-grounded envelope morphing
Experimental setup. Twenty naive listeners were recruited via Prolific.co and participated in the study online, following [institution’s] IRB-approved protocols. Participants saw visual and verbal instructions, including examples of morphing with familiar visual stimuli (faces and shapes) [31] to establish an understanding of a morph as a blend exhibiting features of both inputs.
Stimuli. The auditory stimuli consisted of sequences of 150-ms, 440-Hz pure tones added to a 6-second sample of low-level Gaussian noise. Across trials, we systematically varied one of the three temporal properties described above while holding the others constant. The number of tones (events) ranged from 4 to 16, the inter-onset interval (IOI) varied between 300 and 750 ms, and the onset of the first tone was varied from 10 ms to 3.6 s. Each trial presented two distinct input sounds sequentially, followed by a proposed morphed sound. The Midpoint morph had the linear average of the relevant temporal parameter. The hybrid morphs were temporally intermediate between A and B but were very similar to one of the inputs. The sequential mixdowns were created by sequentially placing the first and second inputs in one audio file. Participants rated appropriateness on a 6-point Likert scale (0 = extremely inappropriate, 5 = extremely appropriate).
Results. Table 1 shows the mean appropriateness ratings (and standard deviations) for a selection of trials illustrating each of the three temporal principles and comparing Midpoint morphs to Unbalanced and Sequence alternatives In all these conditions, the Midpoint morphs received higher ratings compared to both the Unbalanced and Sequence approaches. (Significant differences in a paired-sample t-test are indicated by asterisks).
Table 1: Morph appropriateness ratings for different types of morphs on temporal principles. Asterisks indicate that the rating is significantly lower than that of the Midpoint morph. (*p $<$ 0.05; **p $<$ 0.01; ***p $<$ 0.001)
@@ 4.1 Unsupervised representation learning for envelopes
To morph envelopes, we first train an autoencoder [23] to learn compact envelop representations, providing a bridge to standard morphing methods in deep learning based on interpolation in latent space. We extract amplitude envelopes from the AudioCaps dataset [34], a large corpus of 50k $10$s audio clips covering diverse everyday sounds. To extract the temporal envelopes from the audio files, we compute the root mean square (RMS) envelope using the Hilbert transform [35]. This is followed by lowpass filtering with a cutoff frequency of 30 Hz to remove rapid fluctuations and preserve perceptually salient envelope dynamics as shown in Figure 2. This process yields a large-scale dataset of natural audio envelopes.
@@ 4.2 Supervised learning for perceptually grounded morphing
While the autoencoder (described in Section 4.1) learns compact latent representations, we observe that simply extracting the embeddings of two audio envelopes and interpolating between them results in perceptually unintuitive morphs, similar to those obtained with mixdowns. To address the limitations of naive latent interpolation, we propose a supervised framework that learns to predict perceptually appropriate envelope morphs given a pair of input envelopes and an interpolation weight $\alpha$.
@@
Morphs are constructed by systematically varying one parameter at a time in each set, while keeping others fixed. This procedure generates morphing sequences for impulse count changes, temporal position shifts, density changes and amplitude variations. Invalid configurations (e.g., overlapping impulses) are filtered using automated checks. This synthetic dataset enables supervised training with ground-truth morphs based on perceptual information. Our goal is to build an envelope morphing method that can yield intermediate envelopes controlled by $\alpha$, i.e., ${\texttt{Morph}:E_{a},E_{b},\alpha\mapsto\hat{E}_{\text{morph}}}$. It is likely that human perception is non-linear with respect to $\alpha$. Perceptual mappings of the form ${\texttt{Map}:[0,1]\to[0,1]}$ derived from listening can be incorporated into our method posthoc via ${\texttt{Morph}(E_{a},E_{b},\texttt{Map}(\alpha))}$.
Our morphing model is a twin neural network [24] trained to predict a morphed embedding from a pair of input envelopes and the interpolation weight $\alpha$. The architecture, illustrated in Figure 1, consists of a shared pretrained autoencoder encoder (details in Section 4.1) that maps each envelope to a latent embedding in $\mathbb{R}^{64}$. These two embeddings are then combined using an order-invariant scheme: we compute their sum, their absolute difference, and a weighted average according to $\alpha$. We concatenate these representations to form a feature vector which is passed through a simple three-layer MLP network with a hidden dimension 128. The output is a morphed embedding in $\mathbb{R}^{64}$ which is decoded via the pretrained autoencoder decoder into an output envelope. The model is trained with RMSE loss between the predicted and the ground-truth morph embeddings for 10 epochs with batch size 64.
@@ 5.3 Evaluating the generalization to naturalistic signals
While synthetic data offers interpretability and controlled supervision, it lacks the acoustic complexity and variability of real-world audio. However, fully natural morphing benchmarks lack ground-truth morphs. To balance these needs, we design a naturalistic audio envelope benchmark using real world impulse sounds arranged with synthetic structure [39] as shown in Figure 2. We extract diverse impulse events (e.g., a dog bark, a bicycle bell etc.) from the Clotho dataset [40], manually isolate single impulses, and use them to generate a set of 10k tuples following the same perceptual morphing rules as described in Section 4.2). These audio clips are processed into envelope representations using the same pipeline as before.
@@ 6 Conclusion
We introduced a perceptually grounded framework for morphing temporal envelopes, informed by empirical studies on human auditory perception. By identifying key perceptual factors-such as the quantity, spacing, and placement of impulses-we codified the principles that underlie natural-sounding morphs. These principles enabled the creation of a large synthetic dataset of envelope morphing examples, which we used to train a twin network for generating perceptually valid interpolations. Our experiments demonstrate that the proposed framework consistently outperforms conventional methods across multiple benchmarks. Increasingly, modern sound synthesis and generative models are conditioned on temporal envelopes [41, 42, 43, 44]. By providing a perceptually intuitive way to morph these envelopes, our framework can enhance the controllability and naturalness of the resulting sound morphs in such systems. By bridging psycho-acoustics and DL, this work advances audio generation systems toward perceptually grounded morphing - a critical capability with applications in assistive sound design, immersive media, and computational hearing research.
@@ 7 Acknowledgment
We thank Sony for their support of this work through the Sony Research Award Program (RAP).

### v2

@@ Abstract
Temporal envelope morphing, the process of interpolating between the amplitude dynamics of two audio signals, is an emerging problem in generative audio systems that lacks sufficient perceptual grounding. Morphing of temporal envelopes in a perceptually intuitive manner should enable new methods for sound blending in creative media and for probing perceptual organization in psychoacoustics. However, existing audio morphing techniques often fail to produce intermediate temporal envelopes when input sounds have distinct temporal structures; many morphers effectively overlay both temporal structures, leading to perceptually unnatural results. In this paper, we introduce a novel workflow for learning envelope morphing with perceptual guidance: we first derive perceptually grounded morphing principles through human listening studies, then synthesize large-scale datasets encoding these principles, and finally train machine learning models to create perceptually intermediate morphs. Specifically, we present: (1) perceptual principles that guide envelope morphing, derived from our listening studies, (2) a supervised framework to learn these principles, (3) an autoencoder that learns to compress temporal envelope structures into latent representations, and (4) benchmarks for evaluating audio envelope morphs, using both synthetic and naturalistic data, and show that our approach outperforms existing methods in producing temporally intermediate morphs. All code, models, and checkpoints are available at https://github.com/TemporalMorphing/EnvelopeMorphing.
@@ 1 Introduction
Figure 1: Method overview. In Stage 1, we repurpose an audio autoencoder to learn a latent representation of temporal envelopes. In Stage 2, we train a twin (order-invariant) mapper network (2b) in a supervised setup on synthetic training data that encodes perceptually-grounded morphing rules (2a). During training, the autoencoder is kept frozen while the mapper network learns to morph the embeddings of the illustrated training envelopes.
@@ 3 Perceptually-grounded envelope morphing
Experimental setup. Twenty (10 male, 10 female) US-based listeners were recruited via Prolific.co for an online study, following IRB-approved protocols. They were 20 to 59 years old ($M=39,SD=11.87$). 14 participants identified as White, 5 as Black, and 1 as Asian. 12 participants had 1 to 12 years of musical training ($M=6,SD=6.73$). Participants read instructions with visual examples of morphing (see doi.org/10.1184/R1/29525948) to establish an understanding of a morph as a blend exhibiting features of both inputs [31].
Stimuli. The auditory stimuli consisted of sequences of 150-ms, 440-Hz pure tones added to a 6-second sample of low-level Gaussian noise. Across trials, we systematically varied one of the three temporal properties described above while holding the others constant. The number of tones (events) ranged from 4 to 16, the inter-onset interval (IOI) varied between 300 and 750 ms, and the onset of the first tone was varied from 10 ms to 3.6 s. Each trial presented two distinct input sounds sequentially, followed by a proposed morphed sound. The Midpoint morph had the linear average of the relevant temporal parameter. The hybrid morphs were temporally intermediate between A and B but were very similar to one of the inputs. The sequential mixdowns were created by sequentially placing the first and second inputs in one audio file. Participants rated morphs on a 6-point Likert scale (0 = extremely inappropriate, 5 = extremely appropriate).
Results. The participants showed good inter-rater reliability of appropriateness ratings ($ICC3k=0.77$, $p<0.001$). Table 1 shows the mean appropriateness ratings (and standard deviations) for a selection of trials illustrating each of the three temporal principles and comparing Midpoint morphs to Unbalanced and Sequence alternatives In all these conditions, the Midpoint morphs received higher ratings compared to both the Unbalanced and Sequence approaches (Asterisks indicate significant differences in a paired-sample t-test).
Table 1: Morph appropriateness ratings for different types of morphs on temporal principles. Asterisks indicate that the rating is significantly lower than that of the Midpoint morph. (p $<$ 0.1; *p $<$ 0.05; ***p $<$ 0.001)
@@ 4.1 Unsupervised representation learning for envelopes
To morph envelopes, we first train an autoencoder [23] to learn compact envelope representations, providing a bridge to standard morphing methods in deep learning based on interpolation in latent space. We extract amplitude envelopes from the AudioCaps dataset [34], a large corpus of 50k $10$s audio clips covering diverse everyday sounds. To extract the temporal envelopes from the audio files, we compute the root mean square (RMS) envelope using the Hilbert transform [35]. This is followed by lowpass filtering with a cutoff frequency of 30 Hz to remove rapid fluctuations and preserve perceptually salient envelope dynamics as shown in Figure 2. This process yields a large-scale dataset of natural audio envelopes.
@@ 4.2 Supervised learning for perceptually grounded morphing
While the autoencoder (described in Section 4.1) learns compact latent representations, we observe that simply extracting the embeddings of two audio envelopes and interpolating between them results in perceptually unintuitive morphs, similar to those obtained with mixdowns. To address the limitations of naive latent interpolation, we propose a supervised framework that learns to predict appropriate envelope morphs based on the perception based principles given a pair of input envelopes and an interpolation weight $\alpha$.
@@
Morphs are constructed by systematically varying one parameter at a time in each set, while keeping others fixed. This procedure generates morphing sequences for impulse count changes, temporal position shifts, density changes and amplitude variations. This synthetic dataset enables supervised training with ground-truth morphs based on perceptual information. Our goal is to build an envelope morphing method that can yield intermediate envelopes controlled by $\alpha$, i.e., ${\texttt{Morph}:E_{a},E_{b},\alpha\mapsto\hat{E}_{\text{morph}}}$. It is likely that human perception is non-linear with respect to $\alpha$. Perceptual mappings of the form ${\texttt{Map}:[0,1]\to[0,1]}$ derived from listening can be incorporated into our method posthoc via ${\texttt{Morph}(E_{a},E_{b},\texttt{Map}(\alpha))}$.
Our morphing model is a twin neural network [24] trained to predict a morphed embedding from a pair of input envelopes and $\alpha\in\left\{0,0.25,0.5,0.75,1\right\}$. The architecture, illustrated in Figure 1, consists of a shared pretrained autoencoder encoder (details in Section 4.1) that maps each envelope to a latent embedding in $\mathbb{R}^{64}$. These two embeddings are then combined using an order-invariant scheme: we compute their sum, absolute difference, and a weighted average according to $\alpha$. We concatenate these representations to form a feature vector which is passed through a simple three-layer MLP network. The output is a morphed embedding in $\mathbb{R}^{64}$ which is decoded via the pretrained autoencoder decoder into an output envelope. The model is trained with RMSE loss between the predicted and the ground-truth morph embeddings for 10 epochs with batch size 64.
@@ 5.3 Evaluating the generalization to naturalistic signals
While synthetic data offers interpretability and controlled supervision, it lacks the acoustic complexity and variability of real-world audio. However, fully natural morphing benchmarks lack ground-truth morphs. To balance these needs, we design a naturalistic audio envelope benchmark using real world impulse sounds arranged with synthetic structure [39] as shown in Figure 2. We extract ten diverse impulse events (e.g., a dog bark, a bicycle bell, etc.) from the Clotho dataset [40], manually isolate single impulses, and use them to generate a set of 10k tuples following the same perceptual morphing rules as described in Section 4.2). These audio clips are processed into envelope representations using the same pipeline as before.
@@ 6 Conclusion
We introduced a perceptually grounded framework for morphing temporal envelopes, informed by empirical studies on human auditory perception. By identifying key perceptual factors-such as the quantity, spacing, and placement of impulses-we codified the principles that underlie natural-sounding morphs. These principles enabled the creation of a large synthetic dataset of envelope morphing examples, which we used to train a twin network for generating perceptually valid interpolations. Our experiments demonstrate that the proposed framework consistently outperforms conventional methods across multiple benchmarks. Increasingly, modern sound synthesis and generative models are conditioned on temporal envelopes [41, 42, 43, 44]. By providing a perceptually intuitive way to morph these envelopes, our framework can enhance the controllability and naturalness of the resulting sound morphs in such systems. By bridging psychoacoustics and DL, this work advances audio generation systems toward perceptually grounded morphing - a critical capability with applications in assistive sound design, immersive media, and computational hearing research.
@@ 7 Acknowledgment
Work supported by the Sony Research Award Program (RAP).